OpenStudy (toxicsugar22):

Andy drives 300 miles to Lake Tahoe at 70 miles per hour and returns home at 50 mph. What is his average speed for the round trip. (It is not 6o miles per hour.) Write expressions for the time it takes for each leg of the trip if Andy drives a distance d, at speed r base 1 and returns at speed r base 2

OpenStudy (anonymous):

Let the distance (of one way) be D the time taken to travel at 70 mph = D/70 the time taken to travel at 50 mph = D/50 total time taken = D/70 + D/50 \[{D \over 70} + {D \over 50} = {50D + 70D \over (70)(50)} = {120D \over 3500} = {12D \over 350}\]Total distance covered = D + D = 2D So, avg speed = \[{2D \over { 120D / (70*50)}} = {3500*2 \over 120} = 58.33 mph\]

OpenStudy (anonymous):

did you understand my process ?

OpenStudy (toxicsugar22):

yes

OpenStudy (anonymous):

There's another way to approach this problem which is by calculating the time spent in each trip and divide by the total distance traveled 300/70=x 300/50=y 600/(x+y) = 58.33

OpenStudy (toxicsugar22):

ok so what is the ime trip to Taoe

OpenStudy (toxicsugar22):

time

OpenStudy (anonymous):

the time trip TO taoe would be 300/70 = 4.29 hours

OpenStudy (anonymous):

so did yu understand everthing ?

OpenStudy (toxicsugar22):

ok and the time from trip home

OpenStudy (anonymous):

just like the way there: 300/50 = ?

OpenStudy (toxicsugar22):

ok i got it thanks